# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v2

@@ 4.2. RQ2: What Types of Communities Have Rules Governing AI Use?
To identify subreddit characteristics that suggest the existence of AI rules, we first perform a difference-in-means test between subreddits with and without AI rules. Table 4 presents the results of the test, showing various subreddit characteristics that differ between these two subreddit types. For numerical subreddit characteristics like Subscriber Count and Age the table shows the average value across these two groups of subreddits. For binary characteristics, such as whether a subeddit was created after the launch of ChatGPT or if it has a specific Community Archetype or Topic label, we show the percentage of subreddits in each group with this characteristic. To test for significance we use chi-squared tests for the binary characteristics and the non-parametric Mann-Whitney U test for numerical characteristics. We implement Bonferroni correction to adjust for multiple comparisons.
Table 4 exposes several key differences between subreddits with and without AI rules. Subreddits with AI rules tend to have more subscribers. These subreddits also tend to be newer; correspondingly, a higher percentage of subreddits with AI rules were created after ChatGPT was launched. We observe differences between the groups with respect to their association with the different Community Archetype labels. For example, there is a large and significant difference related to the Content Generation label; subreddits with this label are about $22$ percentage points more likely to have AI rules. In contrast, subreddits of the Social Support Community Archetype are about $6$ percentage points less likely to have AI rules. Lastly, the Art and Celebrity topics are more common in subreddits with AI rules, encompassing about a third of all subreddits with AI rules compared to only $6\%$ of subreddits with no AI rules.
@@
We run several logistic regression (LR) models to validate our results. As a baseline, we repeat the analysis done in Fiesler et al. (2018) and use a LR model to predict the presence of any set of rules for subeddits in the Broad Subreddit Set, using as input variables the following metadata gathered in our second crawl: Subscriber Count [^6] We log transform Subscriber Count to account for the variable’s skewed distribution., Subreddit Age (in years), and a binary indicator of whether or not the subreddit was created after the launch of ChatGPT. We compare these results with a LR model that specifically predicts the presence of AI rules, using as input all of the subreddit characteristics shown in Table 4. Comparing the two regression results isolates the subreddit characteristics that are specifically predictive of AI rules, not rules in general. We find that the Art and Celebrity Topic labels are both predictive of AI rules ($\beta=1.47,p<0.01$ and $\beta=1.41,p<0.01$), supporting the analysis above. In terms of Community Archetype labels, the Content Generation label is predictive of AI rules ($0.010.83,p<$ 0.01). These regression results are in line with the findings from our difference-in-means test.
@@
One possible explanation for why Content Generation communities are more likely to enact rules about AI could be because this type of community is generally less personal, with looser ties between members, and thus may need to rely on formal rules to maintain content quality and standards. Communities that prioritize relationships among users, on the other hand, may be able to use more personal methods to shape user behavior, such as informal norms. For example, consider a Wes Anderson subreddit whose purpose is to discuss “photos that accidentally share Wes Anderson’s unique style.” We labeled this subreddit with both the Art topic and the Content Generation community archetype. The language of this subreddit’s AI rules reflects a distinct concern with the veracity of posted images, rather than concerns about authentic interactions. The subreddit’s rules state that “Your photos don’t have to be images taken by you, but they must be real photos.” Similarly, subreddits labeled with the Celebrity topic tend to group AI rules with bans of photoshop, white borders, or low-resolution images, all of which are concerned with the format and quality of the subreddit’s content. Some Celebrity subreddits seem to encourage AI, but only in cases where it may produce enhanced images. In contrast, subreddits that we labeled with the Social Support community archetype like r/introverts – a subreddit that does not have an AI rule – emphasize in their rules that they care about “authentic connections.” We further explore the implications of these observations in Section 5.
@@ 4.3. RQ3: What Types of Rules Exist for Governing AI?
While our previous analyses explored what types of communities are likely to explicitly regulate AI, we also consider differences in how communities regulate AI. Drawing on the AI Rules Set ($N=4,458$), we produce frequencies of the different Rule Type labels described in Sections 3.3.2 and 3.3.3. Table 5 shows the Rule Type labels used by Fiesler et al. (2018), while Table 6 presents the new labels that we produced via an iterative, inductive coding process. The tables show the percentage of rules in the set that were tagged with each label.
@@
Comparing our results to those of Fiesler et al. (Fiesler et al., 2018) illuminates some interesting differences between AI rules and general subreddit rules. Fiesler et al.’s results found that with general rules, Prescriptive rules, or rules that tell users what to do, occurred in about equal numbers as Restrictive rules that tell users what not to do [^7] These labels are not mutually exclusive: a rule could be tagged as both Prescriptive and Restrictive if it contains elements of both style.. However, in our AI Rules Set, $96.84\%$ of rules were tagged as Restrictive while only $67.7\%$ were tagged as Prescriptive. A Restrictive rule might include statements like “No AI (chatGPT etc.)” or “No AI generated content” while a Prescriptive rule might instead tell users whether to cite if AI is used. This pattern further bears out in the data for our AI Rule: Stance label category, which shows Unqualified Bans (e.g., “No AI-Generated Content”) as the most common label ($55.23\%$ of AI rules). It seems that many subreddits may find it easier to simply ban the use of AI outright, rather than trying to develop policies that can accommodate it under certain circumstances (e.g., requiring the disclosure of AI use). Still, we do see evidence of some communities adopting stances other than outright bans, like the $24.23\%$ of AI rules labeled Stance: Qualified Ban and the $17.95\%$ labeled Stance: Disclosure. Perhaps this implies that some communities will be flexible towards figuring out how to work with, rather than against, this new technology.
@@
Overall, we observe that a high number of AI rules are related to images. Table 5 shows that $46.55\%$ of rules were tagged with the Images label from Fiesler et al. (2018), which according to our codebook, indicates that the rule “is about posting visual content like images, photos, or paintings”. For comparison, Fiesler et al. (2018) tagged less than $8\%$ of the rules in their dataset with this label. We see slightly smaller numbers with our inductively coded labels, which used a codebook that looked for more specific language: Table 6 indicates that $31.4\%$ of AI rules are labeled as Medium: Images and $30.26\%$ are labeled Medium: Art. This observation further supports the finding in Section 4.2 that subreddits with Art and Celebrity topics are more likely to have AI rules, as these are usually image-based subreddits.
@@
Tables 5 and  6 also demonstrate that many AI rules on Reddit focus on concerns about content quality. Around $28\%$ of rules were labeled as relating to Low quality content; $13\%$ and $11\%$ respectively were labeled as Rationale: Low quality and Rationale: Low effort, meaning that they explicitly stated these concerns as their rationale for regulating AI use. Rules about Low quality content emphasize the importance of “creative merit” and content that takes more than “only a few minutes of effort.” These rationales are more prevalent than rationales based on spam, ethics, or accuracy. The focus on quality may suggest that explicit AI regulation is driven by practical concerns about content and reputation, not ideological objections over the use of AI more broadly. This observation echoes the “content quality concerns” shared by moderators in Lloyd et al. (2023).
Table 5. Prevalence of Fiesler et al.’s Rule Type labels in the AI Rules Set ($N=4,458$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules in this set.
@@
Table 6. Prevalence of AI Rule Type labels in the AI Rules Set ($N=4,458$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules included in this set. Examples are drawn from the data but have been shortened for brevity. See the tagging prompt in Appendix A.2 for label definitions.
Rule Category Rule Type % Of Rules Example
@@
Since the Art and Celebrity Topics featured prominently in Section 4.2, we also separately examined patterns among Rule Types within each of these Topics. Among subreddits in the Art Topic, we find a higher share of AI rules with the Copyright and Piracy label ($25.54\%$). These Copyright and Piracy rules primarily focus on proper attribution and posting original content. Among Celebrity subreddits, we see that the Low Quality Content and Format Rule Type labels are much more common ($58.32\%$ and $55.32\%$). The Format label means that references to AI often co-occur with regulations about the format of posts — such as statements about how posts should be titled or specifications about minimum word counts. In terms of the explicit rationales given for the AI rules, almost half of all AI rules in Celebrity subreddits were labeled Rationale: Inauthentic and $34.79\%$ were labeled Rationale: Low Quality. Additionally, $77\%$ of AI rules from Celebrity subreddits explicitly govern the use of “Fakes” (labeled with Medium: Fakes), which upon closer examination appears to be referring to heavily-edited photos (these rules separately make references to “Deepfakes”, which we distinguish with a different label).
@@ 5. Discussion
Our quantitative study offers three main research contributions: (1) a large-scale analysis of rules about AI use in Reddit communities, (2) a taxonomy for classifying these rules, and (3) a dataset of subreddits, their metadata, and public community rules at two distinct points in time. Our study is an effort to support community self-moderation inspired by Seering’s second “guiding question” for the HCI research community: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” (Seering, 2020). The sudden arrival of generative AI poses unique challenges for online communities (Lloyd et al., 2023) that HCI research is uniquely positioned to help address. Our study attempts to measure at scale the prevalence and variety of rules that online communities have enacted in response to this new technology. We hope that our data and analysis can add additional context to qualitative studies into online communities’ responses to AIGC, as well as to general knowledge about online community self-moderation practices. Below we put our findings in dialogue with prior work and discuss design implications for social platforms.
@@ 5.1. The Role of Community Rules
While we focus our analysis on rules about AI use, one of our primary contributions is a dataset of community rules that the HCI research community can use to answer a wider range of questions about community self-moderation. As far as we know, this is the first dataset of its kind, as prior studies of self-governing online communities’ rules (Fiesler et al., 2018; Nicholson et al., 2023; Cai et al., 2021; Hwang and Shaw, 2022; Reddy and Chandrasekharan, 2023) have not made their rules datasets public. Large-scale, quantitative analysis of these rules offers a view into platform-wide patterns that complement past qualitative HCI research into online community self-moderation. Our AIGC-focused analysis finds that subreddit communities, especially those with more subscribers, are changing their rules to respond to generative AI. But what do these rules tell us about the experience of participating in or moderating these communities?
@@
Past research into community self-moderation has found that community rule changes are most often enacted, “in response to unexpected incidents” that impact a community (Seering et al., 2019). Qualitative work on the experience of moderating AIGC (Lloyd et al., 2023) provides evidence that AI rules are created for the same reason, as many moderators report problematic AI use in their communities as a precursor to their formal rules. This mechanism of rule creation suggests that a community’s rules are not a perfect proxy for its values, as community values do not need to be codified into rules unless there is some incident that surfaces community misalignment. Seen in this light, the communities with AI rules in our dataset are not necessarily the communities with the strongest stances on AI, but those where some misalignment about AI use emerged between community members. Viewing rules as evidence of past misalignment could explain why Social Support subreddits were negatively associated with AI rules in our analysis: perhaps these communities are more aligned in general, so that even if they have strong stances on AIGC, they do not need rules to enforce them.
@@
It is worth noting that while rules may indicate some past community misalignment, they do not necessarily tell us how rules shape behavior in practice. For example, prior work has shown that humans cannot reliably detect AIGC (Jakesch et al., 2023), which raises the question of how enforceable AIGC rules are. At the same time, it is unclear whether or not AIGC rules even need to be enforced: past work on community rules suggests that they can shape behavior by clarifying norms (Matias, 2019; Lloyd et al., 2023), even without an effective enforcement mechanism. Still, it is reasonable to assume that some communities with AI rules are attempting to enforce them. We see the prevalence of AI rules, especially in the largest subreddits, as evidence that HCI research should explore design solutions that can help with their enforcement.
@@
One way that HCI research can empower communities to effectively enforce their AIGC policies is by offering tools for policy design. While some online communities are enacting top-down AI policies (Overflow, [n. d.]), Reddit has so far left this decision to individual communities. We like this approach, as past research has shown that community moderators learn by doing and interacting closely with their community (Cullen and Kairam, 2022), resulting in a quick feedback loop that better positions them to craft context-sensitive policies (Seering, 2020). Still, opportunities exist for HCI research to help communities enact effective policies, both by offering empirical evidence of the effects of certain types of rules, as well as new systems for policy development (Zhang et al., 2020). Future work could use our rules dataset to find similar communities that differ in their policy choices, which could be used as a natural experiment to study the effects of different rules. Studying the effects of policies may uncover general knowledge about which kinds of rules (such as “prescriptive” vs “restrictive”) encourage community engagement and which cause members to leave for other communities. Platforms could use this knowledge to offer moderators a reasonable set of “default” rules at community creation time, based on the type of community being created. Some alternative systems for policy development from past HCI research have explored ways of engaging more members of a community, beyond just moderators, in the policy creation process (Zhang et al., 2020). We see these more democratic systems of rule development as promising at a time when norms around AI use are still in flux.

### v3

@@ 4.2.2. Rules Subreddit Set Analysis
To determine which characteristics of subreddits with rules predict the presence of AI rules, we analyze two groups of subreddits in the Rules Subreddit Set: subreddits with rules about the use of AI, and those without such rules. Across these two groups, we consider the presence of the various topic and Community Archetype labels assigned to the subreddits, and implement Bonferroni correction to adjust for multiple comparisons. We first perform difference-in-proportions tests, using the Chi-squared test for significance. Table 4 presents the results of this test. Notice that the topic labels are mutually exclusive, but Community Archetype labels are not. We find that there is a large and significant difference between subreddits with the Content Generation label; subreddits with AI rules are about $22$ percentage points more likely to have this label. In contrast, subreddits with AI rules are about $6$ percentage points less likely to have the Social Support Community Archetype label. There are also significant differences among the topic labels: out of all the mutually exclusive labels, Art and Celebrity are more common in subreddits with AI rules, encompassing about a third of all subreddits with AI rules compared to only $6\%$ of subreddits with no AI rules.
@@
To validate these results we additionally run a logistic regression model. We use all subreddit topic and Community Archetype labels, a subset of which are shown in Table 4, as input variables and use the presence of AI rules as the dependent variable. We get results in line with the findings from our difference-in-proportions tests. For example, the topic labels most predictive of AI rules are Art and Celebrity ($\beta=1.4,p<0.01$ and $\beta=1.69,p<0.01$), supporting the analysis above. In terms of Community Archetype labels, the label that is most predictive of AI rules is Content Generation ($\beta=0.96,p<0.01$), which also supports the above analysis.
@@
Some illustrative examples provide additional context for interpreting these statistical observations. For example, consider a Wes Anderson subreddit whose purpose is to discuss “photos that accidentally share Wes Anderson’s unique style.” Our labeling procedure generated both the Art topic and the Content Generation Community Archetype for this subreddit. The language of this subreddit’s AI rules reflects a distinct concern with the veracity of posted images, rather than concerns about authentic interactions. The subreddit’s rules state that “Your photos don’t have to be images taken by you, but they must be real photos.” Similarly, subreddits labeled with the Celebrity topic tend to group AI rules with bans of photoshop, white borders, or low-resolution images, all of which are concerned with the format and quality of the subreddit’s content. In contrast, subreddits that we labeled with the Social Support community archetype like r/introverts – a subreddit that does not have an AI rule – emphasize in their rules that they care about “authentic connections.” We further explore the implications of these observations in Section 5.
@@ 4.3. RQ3: What Types of Rules Exist for Governing AI?
While our previous analyses explored what types of communities are likely to explicitly regulate AI, we also consider differences in how communities regulate AI. We examine the frequencies of the labels described in Sections 3.3.2 and 3.3.3 for rules in the AI Rules Set ($N=4,458$). Table 5 shows the frequency of rules labeled with Fiesler et al. (2018)’s Rule Type labels, while Table 6 presents the frequency of rules labeled with the AI Rules Taxonomy produced in Section 3.3.2. The tables show the percentage of rules in the set that were tagged with each label.
@@
Comparing the results in Table 5 to a similar analyses of the prevalence of the same labels in a general subreddit rules dataset—performed by Fiesler et al. (Fiesler et al., 2018)—illuminates some interesting differences between AI rules and general subreddit rules. Fiesler et al.’s analysis of general subreddit rules found that Prescriptive rules, or rules that tell users what to do, occurred in about equal numbers as Restrictive rules that tell users what not to do [^7] These labels are not mutually exclusive: a rule could be tagged as both Prescriptive and Restrictive if it contains elements of both style.. However, in our AI Rules Set, $96.84\%$ of rules were tagged as Restrictive while only $67.7\%$ were tagged as Prescriptive. A Restrictive rule might include statements like “No AI (chatGPT etc.)” or “No AI generated content” while a Prescriptive rule might instead tell users whether to cite if AI is used. This pattern further bears out in the data for the AI Rules Taxonomy: the restrictive Unqualified Bans (e.g., “No AI-generated content”) is the most common label ($55.23\%$ of AI rules) in the AI Rule: Stance label category in Table 6. It seems that many subreddits may find it easier to simply ban the use of AI outright, rather than trying to develop policies that can accommodate it under certain circumstances (e.g., requiring the disclosure of AI use). Still, we do see evidence of some communities adopting stances other than outright bans, like the $24.23\%$ of AI rules labeled Stance: Qualified Ban and the $17.95\%$ labeled Stance: Disclosure.
@@
Overall, we observe that a high number of AI rules are related to images. Table 5 shows that $46.55\%$ of rules were tagged with the Images label from Fiesler et al. (2018), which according to the codebook we developed indicates that the rule “is about posting visual content like images, photos, or paintings”. For comparison, Fiesler et al. (2018) tagged less than $8\%$ of the rules in their dataset with this label. We see a similar trend in the labels from our AI Rules Taxonomy: Table 6 indicates that $31.4\%$ of AI rules are labeled as Medium: Images and $30.26\%$ are labeled Medium: Art. The union of these two sets accounted for $57.02\%$ of AI rules in the data. This observation further supports the finding in Section 4.2 that subreddits with Art and Celebrity topics are more likely to have AI rules, as these are usually image-sharing subreddits.
@@
Tables 5 and  6 also demonstrate that many AI rules on Reddit focus on concerns about content quality. In Table 5, around $28\%$ of rules are labeled as relating to Low quality content; in Table 6, $13\%$ and $11\%$ respectively are labeled as Rationale: Low quality and Rationale: Low effort, meaning that these rules explicitly stated these concerns as rationales for regulating AI use. Rules about Low quality content emphasize the importance of “creative merit” and content that takes more than “only a few minutes of effort.” These rationales are more prevalent than rationales based on spam, ethics, or accuracy. The focus on quality may suggest that explicit AI regulation is driven by practical concerns about content and reputation, not ideological objections over the use of AI more broadly. This observation echoes the “content quality concerns” shared by moderators in our earlier qualitative work(Lloyd et al., 2024).
Table 5. Percentage of rules in the AI Rules Set ($N=4,458$) labeled with Fiesler et al.’s non-mututally exclusive Rule Type labels.
@@
Table 6. Percentage of rules in the AI Rules Set ($N=4,458$) with non-mutually exclusive labels from our AI Rules Taxonomy. Examples come from the data but have been shortened for brevity. See the tagging prompt in Appendix A.2 for label definitions.
Rule Category Rule Label % Of Rules Example
@@
Since the Art and Celebrity topics featured prominently in Section 4.2, we also separately examined patterns among Rule Types within each of these Topics. Among subreddits with the Art Topic label, we find a higher share of AI rules with the Copyright and Piracy label ($25.54\%$ vs $16.49\%$ baseline). These Copyright and Piracy rules primarily focus on proper attribution and posting original content. Among Celebrity subreddits, we see that the Low Quality Content and Format Rule Type labels are much more common ($58.32\%$ and $55.32\%$ vs baselines of $27.77\%$ and $19.25\%$). The Format label means that references to AI often co-occur with regulations about the format of posts — such as statements about how posts should be titled or specifications about minimum word counts. In terms of the explicit rationales given for the AI rules, almost half of AI rules in Celebrity subreddits were labeled Rationale: Inauthentic and $34.79\%$ were labeled Rationale: Low Quality, compared with $11.93\%$ and $13.19\%$ respectively in the set of all AI rules. Additionally, $77\%$ of AI rules from Celebrity subreddits explicitly govern the use of “Fakes” (labeled with Medium: Fakes), compared to $15.79\%$ in the wider set. Upon closer examination, rules about “Fakes” in Celebrity subreddits appear to be referring to heavily-edited photos (these rules separately make references to “Deepfakes”, which we distinguish with a different label).
@@ 5. Discussion
Our quantitative study offers three main research contributions: (1) a large-scale analysis of rules about AI use in Reddit communities, (2) a taxonomy for classifying these rules, and (3) a dataset of subreddits, their metadata, and their published community rules at two distinct points in time. Our study aims to inform and support community self-moderation and draws inspiration from Seering’s second “guiding question” for the HCI research community: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” (Seering, 2020). The sudden arrival of generative AI poses unique challenges for online communities (Lloyd et al., 2024) that HCI research is uniquely positioned to help address. Our study attempts to measure at scale the prevalence and variety of rules that online communities have enacted in response to this new technology. We hope that our data and analysis can add additional context to qualitative studies into online communities’ responses to AIGC, as well as to general knowledge about online community self-moderation practices. Below we put our findings in dialogue with prior work and discuss design implications for social platforms.
@@ 5.1. The Role of Community Rules
While we focus our analysis on rules about AI use, one of our primary contributions is a dataset of community rules that the HCI research community can use to answer a wider range of questions about community self-moderation. As far as we know, this is the first dataset of its kind, as prior studies of self-governing online communities’ rules (Fiesler et al., 2018; Nicholson et al., 2023; Cai et al., 2021; Hwang and Shaw, 2022; Reddy and Chandrasekharan, 2023) have not made their rules datasets public. Large-scale, quantitative analysis of these rules offers a view into platform-wide patterns that complement qualitative studies of online community self-moderation. Our AIGC-focused analysis finds that subreddit communities, especially those with more subscribers, are changing their rules to respond to generative AI. But what do these new AI rules tell us about the experience of participating in or moderating these communities?
@@
Past research into community self-moderation has found that community rule changes are most often enacted, “in response to unexpected incidents” that impact a community (Seering et al., 2019). Our earlier, qualitative work on the experience of moderating AIGC (Lloyd et al., 2024) provides evidence that AI rules are created for the same reason, as many moderators report problematic AI use in their communities as a precursor to their formal rules. This mechanism of rule creation suggests that a community’s rules are not a perfect proxy for its values, as community values do not need to be codified into rules unless there is some incident that surfaces community misalignment. Seen in this light, the communities with AI rules in our dataset are not necessarily the communities with the strongest stances on AI, but those where some misalignment about AI use emerged between community members. Viewing rules as evidence of past misalignment could explain why Social Support subreddits were negatively associated with AI rules in our analysis: perhaps the members who participate in these communities are more aligned in their views towards AI, so that explicit rules are not necessary. In general, we suspect that larger, less personal communities, with looser ties between members, may be more likely to rely on formal rules to maintain content quality and standards, as opposed to smaller, closer-knit communities which can use more informal methods, such as community norms.
@@
It is worth noting that while rules may indicate some past community misalignment, they do not necessarily tell us how rules shape behavior in practice. For example, prior work has shown that humans cannot reliably detect AIGC (Jakesch et al., 2023), which raises the question of how enforceable AIGC rules are. At the same time, it is unclear whether or not AIGC rules even need to be enforced: past work on community rules suggests that they can shape behavior by clarifying norms (Matias, 2019), even without an effective enforcement mechanism. Still, our prior work shows that some communities with AI rules are attempting to enforce them (Lloyd et al., 2024). We see the prevalence of AI rules, especially in the largest subreddits, as evidence that HCI research should explore design solutions that can help with their enforcement.
@@
One way that HCI research can empower communities to effectively enforce their AIGC policies is by offering tools for policy design. While some online communities are enacting top-down AI policies (Overflow, 2024), Reddit has so far left this decision to individual communities. Such an approach empowers community moderators who work and interact closely with their community (Cullen and Kairam, 2022), resulting in a quick feedback loop that better positions them to craft context-sensitive policies (Seering, 2020). Still, opportunities exist for HCI research to help communities enact effective policies, both by offering empirical evidence of the effects of certain types of rules, as well as new systems for policy development (Zhang et al., 2020). Future work could use our rules dataset to find similar communities that differ in their policy choices, which could be used as a natural experiment to study the effects of different rules. Studying the effects of policies may uncover general knowledge about which kinds of rules (such as “prescriptive” vs “restrictive”) encourage community engagement and which cause members to leave for other communities. Platforms could use this knowledge to offer moderators a reasonable set of “default” rules at community creation time, based on the type of community being created. Alternative systems for policy development from past HCI research have explored ways of engaging more members of a community, beyond just moderators, in the policy creation process (Zhang et al., 2020). We see these more democratic systems of rule development as promising at a time when norms around AI use are still in flux.